OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = -7 x - 3y + 4z = -14 x - 2y - 3z = -11

OpenStudy (anonymous):

hmmm, math is not my thing but i could do it

OpenStudy (anonymous):

I am just so stuck anything would help

OpenStudy (anonymous):

reduced-row echelon form can solve this. ill explain in a min

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry catmint but i've got nothing!

OpenStudy (anonymous):

thanks for trying Ab590

OpenStudy (anonymous):

no wonder you were stuck...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

what grade are you in, lol?

OpenStudy (campbell_st):

I'd use elimination, the easiest letter to eliminate is y Doing so will leave you 2 equations in 2 unknowns. and then select a letter to eliminate and proceed

OpenStudy (anonymous):

sophomore

OpenStudy (anonymous):

me too, im sophomore!

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

maybe thats why i dun get it

OpenStudy (anonymous):

@campbell_st I dont quite get how you make two equations

OpenStudy (anonymous):

or how to use elimintation

OpenStudy (campbell_st):

ok... the elimination method is exactly that, you work on eliminating a letter from pairs of equations... so you will go from 3 equations with 3 unknowns to 2 equations with 2 unknowns so to eliminate y triple the 1st equation and add it to the 2nd equation 6x + 3y + 3y = -21 x - 3y + 4z = -14 -------------------- 7x +7z = -35 call this equation 4 then you double the 1st equation and add it to the 3rd equation 4x + 2y + 2z = -14 x -2y -3z = - 11 ------------------- 5x - z = -25 call this equation 5 now look at which letter is easiest to eliminate... I'd say z... opposite signs so multiply equation 5 by 7 and add equation 4

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

This is what I have 6x + 3y + 3z = -21 +x - 3y + 4z = -14 -------------------- 7x + 7z = -35 4x + 2y + 2z = -14 +x -2y -3z = - 11 ------------------- 5x - z = -25

OpenStudy (anonymous):

35x= -25/35

OpenStudy (anonymous):

Am I doing this right?

OpenStudy (anonymous):

so confused. if you can tell

OpenStudy (campbell_st):

not quite multiply equation 5 by 7 to eliminate z and add it to equation 4 7x + 7z = -35 35x - 7z = -175 ------------------ 42x = -210 now you can solve for x

OpenStudy (anonymous):

ohhh my thanks. Will you be around to make sure I did it right?

OpenStudy (campbell_st):

well all the working above is correct... you just need to write up the process you used. And you have the solution for 1 variable that you need to include in your answer...

OpenStudy (anonymous):

this is what I wrote. I'd use elimination, the easiest letter to eliminate is Y because it will leave me 2 equations in 2 unknowns. 6x + 3y + 3z = -21 +x - 3y + 4z = -14 -------------------- 7x + 7z = -35 4x + 2y + 2z = -14 +x -2y -3z = - 11 ------------------- 5x - z = -25 7x + 7z = -35 35x - 7z = -175 ------------------ 42x = -210 x= -5

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

thanks for the help